Calculate closing stock from the following details :

Opening stock 4,80,000 ₹
Purchase 13,60,000 ₹
Sales 19,50,000 ₹
G.p. is 30% on cost
1
Expert's answer
2020-09-28T07:59:32-0400

Closing stock"=" Opening stock "+" Purchase"+" Cost of sales.

Cost of sales"=" Opening stock "+" Purchase"-" Closing stock.

Gross profit"=" Sales"-" Cost of sales

Gp"=" "1950000-1840000" "=110000"

Cost of sales"="

"30\over 100" "\times" "110000=33000"

Closing stock"= 480000+1360000+33000=1873000"